OpenStudy (anonymous):

the equation of the line perpendicular to line 3x+2y=6, which passes the intersection point between lines 2x+3y-1=0 and 4x+7y+1=0 will cross Y-axis at point?

OpenStudy (anonymous):

Find the intersection point. Line perpendicular to 3x+2y = 6 is 2x-3y = k. Use intersection point to find k. For Y axis, x = 0. Put x and fnd the required point.

OpenStudy (jack1):

yep, that will work @ShailKumar , but try to give a few more details to make sure they understand all of your steps: 1.Find the intersection point {using simultaneous equations} so equation 1 is 2x+3y-1=0 and equation 2 is 4x+7y+1=0 so use {the substitution method} / whichever method u like to solve for the intersection point. 2.Line perpendicular to 3x+2y = 6 is 2x-3y = k. {because the gradient is always the negative reciprocal} Use intersection point to find k. {by subbing in the values you found in place of "x"and "y". For Y intercept point axis, substitute x = 0 {into the this equation, once you've solved for k: 2x-3y = k } This will allow you to find the y value where the line crosses the Y axis..

OpenStudy (anonymous):

yaa i got it!! thank u :)